• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name A. Camsonne / Z. Ye

    Log entry time 05:44:58 on May 8,2011

    Entry number 358565

    keyword=Trigger issues seems solved

    Zhihong pushed the cables at the TS and that seemed to have fixed the problem. Looking at event type and rates it seems the culprit was T7 one can see it was taking 5 times more triggers than actual scalers. We might want to check/change this cable during the access later in the day.


    FIGURE 1

    FIGURE 2